OpenStudy (luckycharms988):

Solve. 4x - 6 = -3x + 8 X = ?

OpenStudy (anonymous):

U solve it

OpenStudy (luckycharms988):

i only need help.. not the anwser.

OpenStudy (anonymous):

Hey there! We meet once again! :)

OpenStudy (anonymous):

So we need to isolate 'x'

OpenStudy (anonymous):

So, should we add 6 on both sides, or should we subtract 6 on both sides?

OpenStudy (luckycharms988):

Subtract?

OpenStudy (anonymous):

We're subtracting it, so we need to do the opposite!

OpenStudy (luckycharms988):

so add 6 to both sides?

OpenStudy (anonymous):

Yes! So what do you get? :)

OpenStudy (luckycharms988):

10x and on the otherside 11x?

OpenStudy (luckycharms988):

or 4x instead of 10x? because 4x - 6 = -2 + 6 = back to 4

OpenStudy (anonymous):

4x - 6 + 6 = -3x + 8 + 6 4x = -3x + 8 + 6 What is 8 + 6 = ?

OpenStudy (luckycharms988):

14.

OpenStudy (anonymous):

4x = -3x + 14 Now add 3x on both sides 4x + 3x = 14

OpenStudy (luckycharms988):

7x and 17x

OpenStudy (anonymous):

Where did you get 17x? 7x is correct

OpenStudy (luckycharms988):

ok and i got 17x from adding 3x to 14

OpenStudy (anonymous):

You cannot add 3x and 14 since 14 doesn't have any x.

OpenStudy (anonymous):

So now we have 7x = 14 Divide both sides by 7! x = ?

OpenStudy (luckycharms988):

ok i see thx again!

OpenStudy (luckycharms988):

oh ok

OpenStudy (luckycharms988):

14 divided by 2 is 7

OpenStudy (anonymous):

so x = 2!

OpenStudy (luckycharms988):

oh i mean 14 divided by 7 is 2

OpenStudy (luckycharms988):

yes ok thx

OpenStudy (anonymous):

No problem! :)